OpenStudy (anonymous):

find the distance from (-3,1) to 2x-3y=6

OpenStudy (saifoo.khan):

It can be less, it can be more.. it can be anything.

OpenStudy (anonymous):

y=2/3x-2 use points (-3,1) and( 0,-2)

OpenStudy (saifoo.khan):

\[\frac{5}{\sqrt{13}} \approx 4.16025 \]

OpenStudy (anonymous):

'distance' means shortest distance

OpenStudy (anonymous):

\[distance =\sqrt{(-3-0)^{2}+(1-2)^{2}}\]

OpenStudy (anonymous):

sqrt(10)

OpenStudy (anonymous):

is this algebra class or calculus?

OpenStudy (anonymous):

in any case i can write what i did: the point where distance will be shortest is on the line perpendicular to \[2x-3y=6\] through (-3,1) that line is \[3x+2y=-7\]

OpenStudy (anonymous):

the two lines intersect at \[(-\frac{9}{13},-\frac{32}{13})\] and the distance between that point and (-3,1) is not what i wrote! it is what saifoo wrote \[\frac{5}{\sqrt{13}}\]

OpenStudy (anonymous):

which is in fact bigger than \[\sqrt{10}\] so lets see if made a mistake

OpenStudy (anonymous):

oh no i see @ johnny the distance between (-3,1) and (0,-2) is \[\sqrt{18}=3\sqrt{2}\]

OpenStudy (anonymous):

three to the right, three up